OpenStudy (anonymous):

lim ((sqrt(1+h))-1)/h as h goes to 0

OpenStudy (chriss):

The limit is 0

OpenStudy (anonymous):

i got -1/2

OpenStudy (chriss):

Many times you can just plug the "goes to" number into the formula and that will give you the answer. This is one of the cases where that will work. So if you plug 0 in for h, you get \[\sqrt{1-0}-1\] \[\sqrt{1}-1\] 1-1 0

OpenStudy (anonymous):

multiply top and bottom by conjugate (1+h)^(2) + 1

OpenStudy (anonymous):

then try to solve

OpenStudy (chriss):

sorry... the radicand should have been 1+h... but you still get the same result

OpenStudy (anonymous):

u cant just plug 0 into the problem beuase u cant have 0 as the denominator

OpenStudy (anonymous):

((sqrt(1+h))-1)/h ((1+h)^(1/2) - 1)((1+h)^(1/2) + 1)/h((1+h)^(1/2) + 1) = (1 + h - 1)/h((1+h)^(1/2) + 1) = lim 1/((1+h)^(1/2) + 1) x-> 0 1/1+1 = 1/2

OpenStudy (chriss):

right... I misread the problem... I didn't see the denominator your answer is correct

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

where is the negative coming from oh yeah lol didnt see your answer my brain is turned off

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so it is -1/2?

OpenStudy (chriss):

yes, -1/2 is right